Mr. Earl Pearl, accountant for Margie Knall Co., Inc., has prepared the following product-line income data:

Product

Total A B C
  Sales $ 105,000 $ 46,000    $ 25,000    $ 34,000   
  Variable expenses 60,200 29,600    10,800    19,800   
  Contribution margin 44,800 16,400    14,200    14,200   
  Fixed expenses:
    Rent 6,200 2,100    1,800    2,300   
    Depreciation 7,200 2,600    2,000    2,600   
    Utilities 4,560 1,600    660    2,300   
    Supervisors' salaries 5,560 1,100    660    3,800   
    Maintenance 2,780 1,100    760    920   
    Administrative expenses 11,200 2,600    2,800    5,800   
  Total fixed expenses 37,500 11,100    8,680    17,720   
  Net operating income $ 7,300 $ 5,300    $ 5,520    $ (3,520)

The following additional information is available:

The factory rent of $2,300 assigned to Product C is avoidable if the product were dropped.

The company's total depreciation would not be affected by dropping C.

Eliminating Product C will reduce the monthly utility bill from $2,300 to $960.

All supervisors' salaries are avoidable.

If Product C is discontinued, the maintenance department will be able to reduce monthly expenses from $2,780 to $1,700.

Elimination of Product C will make it possible to cut two persons from the administrative staff; their combined salaries total $4,600.

Required:

1. Calculate the advantage or disadvantage in dropping Product C. (Input the amount as a positive value. Omit the "$" sign in your response.)
    (Click to select)  Advantage  Disadvantage  in dropping Product C $   
2. Should the product be dropped?
  • Yes

  • No
Answer #1

1)Product line income data after removing the product C

Total sales =46000+25000= 71000

Variable expenses =29600+10800= 40400

Contribution margin =71000-40400= 30600

fixed expenses

Rent =2100+1800= 3900 (removed the rent expense assigned to product C)

Depreciation =7200 (Company's total depreciation will not be affected if product c is removed)

Utilities =1600+660+960= 3220 (utility expenses will reduce to 960 when product c is dropped)

Supervisor salaries =5560-3800= 1760 (supervisor expenses incurred related to product C is removed)

maintenance expenses =1700 (will reduce as the product c is dropped)

Administrative expense =11200-4600=6600

New Total fixed expenses=3900+7200+3220+1760+1700+6600=24380

New Net operating income=30600-24380=6220

The difference in net operating income as a result of dropping product C =7300-6220=1080

(7300 is the net operating income when product a,b, c are produced)

Here it is found that the net operating income has come down from 7300 to 6220 as a result of dropping product C since some of the fixed expenses still incurred by the factory. (But if all the fixed expenses associated with product C was avoidable then the net operating income would be higher than the first net operating income).

So it is a disadvantage for the company in dropping the product C as it net operating income decreases

2)The product C should not be dropped